OpenStudy (anonymous):

Solve using the addition method. What is the x-value of the solution to the system? 2x – 3y = -1 3x + 4y = 24

OpenStudy (anonymous):

i dont get itt :'((((

OpenStudy (saifoo.khan):

where are u stuck?

OpenStudy (anonymous):

u sed to multiply the first equation by 2 but wat do i do after tht like ill get 4x-6y=-2 ???

OpenStudy (saifoo.khan):

No, in this one, you have to multiply the second one as well.

OpenStudy (saifoo.khan):

first by 3 second my -2

OpenStudy (anonymous):

2x – 3y = -1 6x-9y= -3 3x + 4y = 24 6x -8y= -48

OpenStudy (saifoo.khan):

2x – 3y = -1 6x-9y= -3 3x + 4y = 24 -6x -8y= -48

OpenStudy (anonymous):

-17y=-51

OpenStudy (anonymous):

y= 3?

OpenStudy (saifoo.khan):

Great.

OpenStudy (anonymous):

lozl but how do i get x?

OpenStudy (saifoo.khan):

now insert y = 3 in any equation u like.

OpenStudy (anonymous):

ok into the new one or the original?

OpenStudy (saifoo.khan):

anyone you like, it will give u the same answer.

OpenStudy (anonymous):

6x-9y= -3 6x-9(3)=-3 6x-27=-3 +27 +27 6x=24/6 x=4??

OpenStudy (saifoo.khan):

wooow, great. great/ woot woot.

OpenStudy (anonymous):

woot!!! ty saifoo!!

OpenStudy (saifoo.khan):

you welcome.